OpenStudy (anonymous):

(x+2)(2x+3)=6 My options: (-2, -3/4) (-7/2,0) (0)

OpenStudy (anonymous):

First Expand the lhs to get \[2 x^2+7 x+6=6 \]

OpenStudy (anonymous):

Second Step \[2 x^2+7 x=0 \]

OpenStudy (anonymous):

Third Step \[ x( 2 x +7)=0 \]

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

Wait, how did they get in the parentises?

OpenStudy (anonymous):

You factor the x out

OpenStudy (anonymous):

Im trying to figure it out and I'm going backwards.

OpenStudy (anonymous):

OK. Take your time

OpenStudy (anonymous):

Wouldn't be (-7/2,0) because you can't factor a 7?

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

Yes x=0 x=-7/2

OpenStudy (anonymous):

Ha! Lol thank you!